Bowling ball that is rolling has a mass of 4kg at a speed of 12m/s and reaches a height of 3.5 meters how much work did friction do?


1
Expert's answer
2021-04-26T17:12:16-0400

"v=at,"

"h=\frac{at^2}{2}=\frac{v^2}{2a},"

"a=g-\frac Fm,"

"\frac Fm=g-\frac{v^2}{2h},"

"F=gm-\frac{mv^2}{2h},"

"A=-Fh=\frac{mv^2}{2}-mgh,"

"A=\frac{4\cdot 12^2}{2}-4\cdot 9.8\cdot 3.5=150~J."